### v1

# Exploration of Masked and Causal Language Modelling for Text Generation

Abstract
Large Language Models (LLMs) have revolutionised the field of Natural Language Processing (NLP) and have achieved state-of-the-art performance in practically every task in this field. However, the prevalent approach used in text generation, Causal Language Modelling (CLM), which generates text sequentially from left to right, inherently limits the freedom of the model, which does not decide when and where each token is generated. In contrast, Masked Language Modelling (MLM), primarily used for language understanding tasks, can generate tokens anywhere in the text and any order. This paper conducts an extensive comparison of MLM and CLM approaches for text generation tasks. To do so, we pre-train several language models of comparable sizes on three different datasets, namely 1) medical discharge summaries, 2) movie plot synopses, and 3) authorship verification datasets. To assess the quality of the generations, we first employ quantitative metrics and then perform a qualitative human evaluation to analyse coherence and grammatical correctness. In addition, we evaluate the usefulness of the generated texts by using them in three different downstream tasks: 1) Entity Recognition, 2) Text Classification, and 3) Authorship Verification. The results show that MLM consistently outperforms CLM in text generation across all datasets, with higher quantitative scores and better coherence in the generated text. The study also finds no strong correlation between the quality of the generated text and the performance of the models in the downstream tasks. With this study, we show that MLM for text generation has great potential for future research and provides direction for future studies in this area.
Index Terms– Masked Language Model, Causal Language Model, Text Generation, Synthetic Data Applications
I Introduction
The rise of Large Language Models (LLMs) has gained much attention in recent years. These models have demonstrated the remarkable ability of complex reasoning and generate human-like text. However, it became apparent since its very introduction that the transformer architecture, the backbone of LLMs, has numerous limitations. In fact, the attention mechanism [1], the core component of transformers, struggles with capturing long-range dependencies and does not scale efficiently to capture longer sequences. This finding led researchers to explore various ways to overcome these limitations and introduce many transformer variants [2, 3, 4]. This collective effort also led to the exponential growth in size of these models.
Despite the tremendous effort put into language generation research, the prevalent and "almost unchallenged dogma" in text generation is that models must produce language sequentially, uni-directionally, and from left to right. The core idea is that a model would take an input text or prompt, e.g. see Table I, and then iteratively generate the next token until completion.
Step Generated Text
Prompt The quick brown fox
1 The quick brown fox jumps
2 The quick brown fox jumps over
3 The quick brown fox jumps over the
4 The quick brown fox jumps over the lazy
5 The quick brown fox jumps over the lazy dog
TABLE I: An example of sequential, unidirectional, left-to-right text generation using a CLM model.
This process is called Causal Language Modelling (CLM) because the probability of generating each next token is conditioned only on the tokens that appear before it in the sequence. This unidirectional approach inevitably restricts their capacity to leverage the full context of a text, as these models are prevented from incorporating or generating tokens that appear later in the sequence. This limitation is not present when using another approach called Masked Language Modelling (MLM).
MLMs learn to generate tokens by looking at a bidirectional context. Existing research has almost exclusively focused on using MLM to pre-train models for Natural Language Understanding (NLU) tasks, such as sentiment analysis, text classification, named entity recognition, etc [5]. However, previous research demonstrated that MLM models could be used for text generation [6] and predict tokens by considering context from both directions of a given sequence, as shown in Table II.
Step Generated Text
Prompt The quick brown fox [MASK] [MASK] [MASK] [MASK] [MASK]
1 The quick brown fox [MASK] [MASK] [MASK] [MASK] dog
2 The quick brown fox jumps [MASK] [MASK] [MASK] dog
3 The quick brown fox jumps [MASK] [MASK] lazy dog
4 The quick brown fox jumps [MASK] the lazy dog
5 The quick brown fox jumps over the lazy dog
TABLE II: An example of text generation using MLM.
Despite their potential, MLM applications in text generation, particularly for tasks such as synthetic data generation [7], have been greatly overlooked, leaving their generative capabilities massively unexplored. In this paper, we aim to challenge this assumption in text generation by comparing CLM with MLM using models of comparable size. By exploring this potential, we aim to find new directions for this NLP area and understand under what conditions MLM for text generation can shine. Since there is little research on this topic, our starting point will be synthetic data generation, as we found that this approach would be easier to evaluate since we can easily compare the synthetic dataset with the original one and also use the synthetic data to train models in downstream tasks and evaluate their performance.
We raise the following research questions:
  1. 1.
    Can MLM outperform CLM in text generation, and in what contexts?
  2. 2.
    Does domain-specific knowledge used for pre-training and fine-tuning matter for generating text that is of higher quality with regard to the quantitative metrics?
  3. 3.
    Does higher quality in the texts equal better performance in downstream tasks?
  4. 4.
    What kind of tokens provide enough context to generate high-quality text?
To answer these questions, we aimed to thoroughly evaluate and compare the ability of CLM and MLM models of comparable sizes to generate text. The following are the main tasks of this investigation:
  1. 1.
    Fine-tune a number of Pre-trained Language Models (PLMs), including T5, BART, and BERT and their variations.
  2. 2.
    Explore different generation approaches by masking different tokens and establish which ones yield the best performance.
  3. 3.
    Compare the generations of MLM and CLM models with standard quantitative metrics and qualitatively assess the generations.
  4. 4.
    Examine the usefulness of the generations of the fine-tuned models in downstream tasks, namely Named Entity Recognition (NER), Text Classification and Authorship Verification.
  5. 5.
    Interpret the results of the evaluations and conclude the optimal model types and masking strategies for different text generation tasks across the three different datasets.
To the best of our knowledge, this is the most extensive investigation that involves the comparison of MLM and CLM models for text generation.
II Background and Related Work
II-A Generative Models
In this section, we introduce the generative models that we used in our experiments. These models are different both in their architectures and in the domains they were trained in. This will allow us to more thoroughly compare how their different characteristics affect the quality of the generations.
II-A1 BERT family of models
Bidirectional Encoder Representations from Transformers (BERT) [5]. BERT is designed to pre-train deep bidirectional representations by conditioning on both left and right contexts in all layers. As a result, the pre-trained BERT model can be fine-tuned with just one additional output layer to create state-of-the-art models for a wide range of tasks, such as question answering [8] and natural language inference [9], without substantial task-specific architecture modifications. The model has 345M parameters.
RoBERTa Robustly Optimised BERT Pretraining Approach (RoBERTa) [10] builds upon BERT by modifying key hyper-parameters in its pre-training process, which includes training the model longer, with more data and on bigger batches. This model showed better performance than BERT across a wide range of tasks. The model has 355M parameters.
BiomedNLP-PubMedBERT BiomedNLP-BiomedBERT-large-uncased-abstract (PubMedBERT) [11] is a BERT variant fine-tuned for the biomedical domain that uses a vast corpus of biomedical texts from PubMed, making it more effective in handling the biomedical language. The model showed state-of-the-art performance on the Biomedical Language Understanding and Reasoning Benchmark (BLURB) [^1] https://microsoft.github.io/BLURB/. We want to use this model because we want to understand whether task-specific knowledge helps the model generate higher-quality text for this specific task.
II-A2 T5 family of models
Text-to-Text Transfer Transformer (T5) [12] is an encoder-decoder model that frames all NLP tasks as text-to-text problems. It is pre-trained on the Colossal Clean Crawled Corpus (C4) using a denoising objective, where the model learns to reconstruct corrupted input sequences. Providing task-specific prefixes to the input text allows it to be fine-tuned for various tasks, such as translation, summarization, and question-answering. The t5-large version is made up of 770M parameters, which is a comparable size to the large models in the BERT family.
SciFive-large-Pubmed_PMC SciFive [13] is a version of the T5 model fine-tuned on scientific and biomedical literature, specifically texts taken from PubMed and PMC (PubMed Central). SciFive is excellent in tasks related to scientific text, such as document summarisation, entity extraction, and knowledge graph construction. The model is used primarily for understanding and generating scientific text. Simarly to PubMedBERT (see Section II-A1), we want to use this model to evaluate whether task-specific knowledge affects the quality of the generations.
II-A3 BART
BART (Bidirectional and Auto-Regressive Transformer) [14] is a generalized pre-training model based on the Transformer architecture. The model was pre-trained using 5 techniques:
  • Token masking, in the same way as used in BERT, randomly masks tokens.
  • Sentence permutation shuffles the sentences of a document randomly.
  • Document rotation rotates the sentence so that it begins with a randomly selected token.
  • Token deletion takes the original sentence and randomly deletes a token from the text.
  • Text infilling works in the same way as token masking but masks word sequences instead of single words.
The models has 406M parameters.
II-B Data Augmentation
Data augmentation has recently been used more extensively in NLP due to more work performed in low-resource domains [7] and the increasing popularity of large neural networks that require massive amounts of training data [15]. Applying it to this field is far more challenging than in others due to the discrete nature of language, where slight modifications can significantly alter the semantics of a sentence. Notwithstanding these challenges, numerous works have applied data augmentation for NLP. These can be simple approaches, often just involving rule-based approaches that can offer minor improvements [16]. More complex approaches are often model-based and have been applied to numerous sub-fields of NLP, such as text classification [17], question answering [18], and machine translation [19].
II-C Synthetic Data Generation
In contrast to data augmentation, which relies on bootstrapping existing data, synthetic data generation creates entirely new datasets from scratch. What this method essentially does is generate a fresh new dataset that has the same properties and distribution as the real-world dataset. Generative AI and Synthetic Data Generation started when Variational Autoencoders (VAEs) [20] and Generative Adversarial Networks (GANs) [21] were introduced, demonstrating the ability to generate realistic images, signals and tabular data.
Synthetic data in NLP became necessary as the performance of Language Models was shown to be limited when evaluated on domain-specific benchmarks. ClinicalT5 [22], for example, was a version of T5 that was fine-tuned on the medical corpus. It showed better performance than T5, which was initially just trained on general text data. Most specialised domain datasets are challenging to access, showing the importance of synthetic data becoming a proxy of the original datasets for model fine-tuning [7].
II-D Masked Language Modeling for Text Generation
The idea that BERT [23] could be used for text generation was first introduced in 2019 [6]. This approach generates some or all the tokens simultaneously by looking at a bidirectional context. Non-autoregressive text generation was primarily used in Neural Machine Translation (NMT) to speed up inference time, by decoding multiple tokens simultaneously, at the cost of sacrificed translation accuracy compared to its counterpart, autoregressive generation [24]. However, few works have explored the potential of BERT in generating text outside of machine translation [25, 26]. However these works focused on instruction following and story generations, respectively, but still focused on improving inference speed as the primary goal rather than text quality. In addition they did not use any domain-specific dataset, and did not attempt to utilize the generations in downstream tasks. In our work, we aim to expand on this by examining the conditions under which MLM models can generate better text than CLM models, thereby finding the best direction for future research in MLM for Text Generation.
III Methods and Experimental Design
Figure 1: Evaluation Pipeline
III-A Problem Formulation
Let $C$ be a space of text features representing individual texts, such as single tokens. Let $L$ be a set of unmasked tokens, which is our context for generating synthetic data. Thereby, we have a dataset $D_{\mathcal{L}}^{C}$ containing a set of texts.
For each combination of tokens $\mathcal{L}\subseteq L$, we initially have a subset of data $D_{\mathcal{L}}$ defined as $D_{\mathcal{L}}=\{D_{\mathcal{L}}^{n}\}_{n=1}^{N_{\mathcal{L}}}$ containing texts associated with the token combination $\mathcal{L}$. Individual texts are indexed by $n$ for each $\mathcal{L}$, where $N_{\mathcal{L}}$ is the number of texts for the token combination $\mathcal{L}$.
Our objective is to generate a synthetic dataset of texts that replaces the real dataset entirely, conditioned on the combinations of relevant tokens $\mathcal{L}$. To achieve this, we aim to learn a density function $\hat{d}\{C|\mathcal{L}\}$, approximating the true distribution $d\{C|\mathcal{L}\}$ of the clinical instructions conditioned on each token combination $\mathcal{L}$.
Once the distributions are learned, we generate synthetic sub-datasets by drawing random variables from $\hat{d}\{C|\mathcal{L}\}$ for each $\mathcal{L}$. We then combine the sub-datasets to form our final dataset, which effectively replaces the original dataset while preserving its essential characteristics.
III-B Datasets
III-B1 Medical Dataset
For the medical task, we used the Medical Information Mart for Intensive Care (MIMIC-III) database [27, 28]; in particular, we focused on the section used for the National NLP Clinical Challenges (n2c2) 2018 shared task data on adverse drug events and medication extraction with gold labels [29]. We then extracted the discharge summaries for each clinical note from the dataset. We decided to use this specific dataset because it contains many caregiver notes and medication prescriptions for various clinical conditions and treatments, enabling our model to generate diverse samples.
III-B2 Movive Plot Synopsis Dataset
For the second task, we used the Movie Plot Synopses with Tags (MPST) Corpus [30]; the dataset comprises a corpus of 14K movie plot synopses and is associated with 70 tags per synopsis, retrieved from Wikipedia and IMDb.
III-B3 Authorship Verification Dataset
The PAN Author Verification dataset [^2] https://pan.webis.de/clef23/pan23-web/author-identification.html contains texts from 112 individuals, each providing a written and spoken text of different Discourse Types on various topics and varying degrees of formality. These texts were transcribed into essays, emails, interviews, and speech transcriptions. All the texts in the dataset are from native English speakers.
III-C Training and Generation
III-C1 MLM Models
Training We used the standard MLM training protocol to train the MLM models. This approach entails masking random tokens and then letting the model learn to predict these based on the provided context, i.e., the non-masked tokens. While typically the proportion of masked tokens in the training phase is about 15% [5], in our case, we set the proportion to 50% random tokens to let the model learn in conditions that more closely align with the generation phase.
Generation The process we implemented for generating text using the MLM models consists of two steps. First, the original text would be corrupted according to the specific corruption method described in Section III-C3 to contain masked and unmasked tokens. Second, the model would predict each [MASK] token one by one until no more mask tokens remained in the text. As shown in Figure II, the model does not need to follow a particular order to unmask each token and does so based on which prediction for each [MASK] token has the highest confidence.
III-C2 CLM Models
Training To train the CLM models, we followed a similar approach as for MLM, but this time, we removed all of the masked tokens from the text and input the unmasked tokens into the model, which is then trained to generate text given the list of unmasked tokens.
Generation Generation works analogously to training: we first select the input tokens and then let the model generate the text given that context.
III-C3 Generation Approaches
To select the right tokens to be the input for the generation, we employed different approaches and assessed their impact on the performance of the generations of the models models. We used the same approaches for both MLM and CLM. The only difference was that in MLM, we replaced the tokens with [MASK], while in CLM, we removed the tokens from the input.
The following are the approaches that we used:
  • Random: The tokens were picked randomly from the input text. With this approach, we masked a percentage of tokens from 10% to 100%.  [^3] For CLM models, we retained that the evaluation of the 100% masking ratio was unfair, so we excluded that from the final evaluation. The goal of this approach was to establish the baseline for text generation and to determine the extent to which the ratio of the number of masked tokens can impact the quality of the generated text by the models.
  • Stopwords: This time, we masked all the tokens in the text that were not English stopwords. The goal of this approach was to understand whether punctuation alone could provide enough context to generate high-quality text.
  • Punctuation: Same as above, except that we masked everything except punctuation.
  • Stopwords and Punctuation: we also sought to evaluate the quality of the generations when both stopwords and punctuation were included in the input text.
  • NER: with this approach, we used SciSpacy  [^4] https://allenai.github.io/scispacy/ to detect all of the entities appearing in the text and masked the rest. This approach aimed to understand how the model can vary in its generations while maintaining the correct important entities. This generation approach was applied only to the Medical dataset, given that it was the only one with easily recognisable entities that are the core part of the text.
Figure 2: Experimental Design of the Downstream NER Task.
III-D Downstream Tasks
In addition to the quality of the generations, in the first two tasks we wanted to understand their usability to train other models. In the first two tasks wanted to understand to what extent the quantitative metrics correlated with the usefulness of our texts. With the third task instead, we sought to determine which approach could generate text most resembling the original style according to a third model.
III-D1 Named Entity Recognition (NER) Task
The goal of NER is to locate and classify named entities mentioned in unstructured text into categories such as person names, organizations, locations, etc. In this task, we want to evaluate the performance of an NER model trained on the generated texts. Since the dataset is medical in this case, we will focus on the ability of the model to detect medical entities.
As Figure 2 shows, we first used SciSpacy(a NER  [^5] https://allenai.github.io/scispacy/ to detect medical entities in the synthetic text. The labels generated by SciSpacy would then be fed to the Spacy model (see SectionIII-D1) along with the synthetic text for training. Additionally, we used SciSpacy to extract the test NER labels from the test dataset. We then fed the test dataset to the trained Spacy model and compared its predictions with the NER labels extracted by SciSpacy, computing the F-1 Score.
NER model: SpaCy
III-D2 Text Classification Task
Text Classification is associating labels to a corpus of texts. We perform a multilabel binary classification task. The dataset is from Authorship Verification 2023  [^6] https://pan.webis.de/clef23/pan23-web/author-identification.html. We further processed it to contain only human-generated plot synopses and reduced the number of tags to 6 to maintain only the ones appearing in most samples. This task focuses on the model’s ability to classify texts correctly. Like in the NER task, we trained a BiLSTM model on the generated texts but left the tags unchanged. The trained models were evaluated on the same testing set to measure their movie-level f-1 score, which is computed with the total number of True Positives, True Positives and False Negatives across all the categories.
With this task, we wanted to determine whether the texts generated by our models could be successfully used to train a text classification model, how much text quality mattered for this task, and which specific tokens provided enough context to generate useful text.
Figure 3: Experimental Design of the Downstream Text Classification Task.
The BiLSTM model was trained on the generated plot synopses with their corresponding tags taken from the real datasets. We then used the trained model to process the test dataset and predict the most likely tags. Finally, we compared the predictions with the true test labels and computed the movie-level F-1 score.
III-D3 Authorship Verification Task
Authorship verification is the task of analysing the linguistic patterns of two or more texts to determine whether the same author wrote them. Initially, experts performed this analysis by considering spelling mistakes, grammatical inconsistencies and other linguistic features. However, more recently, machine learning algorithms have been trained to accomplish the same. In this task, we collected all the pairs of texts that our fine-tuned model originally predicted belonged to the same author 4. We then altered one of the texts in each pair, and we tried to determine if the prediction of the model about the authors being the same remained unchanged despite the corruption.
Figure 4: Experimental Design of the Downstream Author Verification Task.
As Figure 4 shows, for this task first, we give 2 a pair of texts to the classification model. If the model outputs "Yes", we then corrupt Text 2 and reconstruct it using our generative model. Finally, we classify the pair again and count the number of times the model correctly outputs "Yes" again. We focused only on when the model outputs "Yes" initially because it is harder for two texts to come from the same author than the opposite.
The third downstream task does not directly involve training models and instead it essentially aims to determine whether the generated text was of high quality enough to "fool" the prediction model into thinking that the text still comes from the same author by maintaining a similar writing style to the original text.
Author Verification Model: Flan-T5-XL Fine-tuned LAnguage Net T5 XL (FLAN-T5-XL) [31] is an extension of the T5 model designed for tasks requiring a deep understanding of language context. It is trained on the "text-to-text" framework (see Section II-A2). The model was developed to address a range of text-based tasks, including translation, summarisation, question answering, and text generation. In this context, the "XL" stands for the size of the model, which is 3 billion parameters in this case. The model has already been trained on the Authorship Verification task.
IV Results and Discussion
In the following section, we first show the results of the quantitative, qualitative, and downstream evaluations. We then discuss the results and their possible implications. Due to computational constraints when generating text with MLM (as explained in Section V-A), the synthetic datasets were between 130 and 200 samples in size.
IV-A Quantitative Evaluation
After computing the scores for the quantitative evaluation, we plotted them in the following figures. The evaluation metrics we used compare the generations obtained from corrupted text and the original texts before they were corrupted. The scores of all of the masking approaches (introduced in Section III-C3) we used are present in the figures; scores from random masking are shown as the masking ratio increases, while scores obtained by the other approaches are added to the graph as flat lines. See Appendix LABEL:app:quantitative_results_graphs for the complete collection of figures for each dataset. The MLM models are RoBERTa, BiomedNLP-PubMedBERT, and BERT, whilst the CLM models are T5, SciFive, and BART.
IV-A1 Medical Dataset
Figure 5: Medical BLEU Scores.
As shown in Figure 5, the MLM model generations (Roberta-large & BiomedNLP-PubMedBERT) align more closely with the original text and, therefore, achieve higher scores. It is clear that MLM models are more sensitive to changes in the masking ratio, and they get better scores when the ratio is lower. They also show very low scores when the ratio is 100%. With regard to CLM models, they show low scores across all generation approaches; in this case, a low BLEU score shows low precision. The punctuation and stopwords masking are also relatively good for the MLM models showing mid-range scores, while in CLM, they are at the bottom of the graph.
Figure 6: Medical ROUGE Scores.
Figure 6 shows that while still worse than the MLM models, the ROUGE scores of the CLM models are relatively better than the BLEU scores, meaning that at least some of the words that appear in their generations also appear in the reference texts (recall).
Figure 7: Medical METEOR Scores.
As shown in Figure 7, the METEOR evaluation, a recall-oriented metric, confirms our hypothesis above, showing relatively good scores for the CLM models compared to the BLEU scores. However, it is clear that CLM models include many tokens in their generations that do not appear in the reference texts compared to the MLM models.
Figure 8: Medical BERT Scores.
In Figure 8, the BERTScore shows that notwithstanding the specific medical language, all models can generate samples that are semantically similar to the reference texts. This shows good adaptability to the task-specific domain for all models, with MLM models showing slightly better performance.
IV-A2 Movie Dataset
Figure 9: Movie BLEU Scores.
Again, as expected, Figure 9 less masking leads to higher scores. When generating samples from the movie dataset, the MLM model still achieves higher scores. While t5-large’s score is as low as with the medical dataset, bart-large improved significantly, achieving good scores overall. In addition, the model seems more sensitive to the level of context, meaning that the model can use it better and produce informed generations. Neither the punctuation nor stopwords approaches provide enough context for the model to generate good-quality samples, with their scores being at the bottom of the graph.
IV-A3 Author Dataset
Figure 10: Author BLEU Scores.
As with the other two datasets, Figure 10 shows that MLM models perform better than the CLM models; this time, bart-large performs well in the variable rations compared to the previous two datasets, while t5-large is not able to produce accurate generations.
IV-B Human Evaluation on Sampled Data
For this task, we picked the first 5 samples generated by the models when using a random 50% masking ratio.
IV-B1 Medical Generations
MLM Generations These samples were generated by Roberta large.
please shower daily including washing incisions gently with mild soap, no baths or swimming pat and dry at all times please no powder, powder, cream, or ointments to cover each incision please weigh yourself and measure yourself every evening and every morning, these should be written down on the chart no driving for one month or while taking narcotics. driving will be discussed at follow up appointment with surgeon please no lifting more than pounds for weeks please call with any questions or concerns telephone or fax please call cardiac surgery office with any questions or concerns telephone or fax.
TABLE III: Roberta Medical Generations based on 50% random context.
As shown in Table III, MLM generations are high in quality. They show few hallucinations, a high level of coherence in the texts, and a good length that matches the expected length for discharge summaries. In addition, these texts show an appropriate structure for medical discharge summaries because they include relevant sections such as patient instructions, medication changes, and follow-up recommendations. What is also remarkable is that the model demonstrates a good grasp of medical terms, such as "angioedema," "pneumonia," and "endoscopy." The texts are generally coherent; however, information is sometimes introduced out of nowhere. For example, in the fourth summary, "mm nodular area on the chest CT" is disconnected from the rest of the summary.
CLM Generations These samples were generated by T5 large.
be at follow up please call with any or concerns/fax cardiac surgery office any or concerns/fax cardiac surgery office any or concerns telephone. answering service contact on hours telephone. answering service contact on hours telephone. answering service contact on hours telephone. answering service contact on hours telephone. answering service contact on hours telephone. answering service contact on hours telephone. answering service contact on hours telephone. answering service contact on hours telephone. answering service contact on hours telephone. answering service contact on hours telephone. answering service contact on hours telephone. answering service contact on hours telephone.’, 1: ’take any new. Please take any new of regular medications take any new as ordered. Please of regular medications take any new. Please of regular medications take any new. Please of regular medications take any new. Please of regular medications take any new of regular medications take any new. Please. Please to to. Please to chest pain to to chest to chest to chest to to……. take….
TABLE IV: T5 Medical Generations based on 50% random context.
In Table IV, the texts show many repetitions; multiple phrases are repeated for almost the entire length of the texts. For example, in the first generated text, the phrase "answering service contact on hours telephone" is repeated multiple times. First of all, the generated texts lack the structure and organization that proper medical discharge summaries would require; second, most of the information appears randomly generated and without a clear flow. This all together makes it very difficult to understand the key instructions given by the summary. Some texts contain incomplete sentences, such as "take any new. Please take any new or regular medications as ordered. Please, of regular medications, take any new." Over longer sentences, it is clear that the CLM model cannot maintain any coherence in the texts; indeed, while the texts include medical terms and phrases, they are often used in a disjointed way and lack context. For instance, "systolic. blood systolic. blood systolic. blood systolic. blood systolic. blood primary care doctor to hospital with a reaction called angioedema." is a clear hallucination. There are many instances of irrelevant information being generated, such as "if it becomes black or blood in stool, go to the hospital." which appears in the summary without appropriate context.
IV-B2 Movie Generations
MLM Generations These samples were generated by Roberta large.
1. The film opens with a montage of scenes of daily life in Australia during the apocalypse: a rural school in the desert, a busy street in a small town, a traffic jam in the city, all being affected by the adverse weather conditions that suddenly appear. Only the local Aboriginal people seem to recognize the changes. The film’s first scene is one of these "freak occurrences" in Sydney, where it occurs among a group of men in a pub, all affected by the mysterious effects of the weather. Arriving at the court of public safety, the case is ruled not guilty; and the men are given the opportunity to use the public legal aid system, where Burton is the only legal counsel procured for their defense. The circumstances under which he was contacted and retained are unusual, in that his law practice is corporate, and not criminal defense. The film focuses on the case, and the professional and personal lives of the unravel involved, including bizarre dreams that Burton relates to the court. Interacting interviewing one of the accused (David Gulpilil), he is also able to uncover a strange ritual that has taken place in his home, and connect his rituals with the murder case, which turns out to have been an Aboriginal ritual performed by curse same shaman that the victim believed. Learning more about Aboriginal practices and the idea of Dreamtime as a parallel world of existence, Burton concludes that this strange ritual bodes ill for the apocalypse. The film ends with a confrontation between Burton and the victim’s shaman at a subterranean burial site. Overcoming the shaman and Burton reach the surface, buried in a tunnel are various tribal relics. After emerging from the tunnel, he sits on the beach and stares out at the horizon. The screen is filled with the shot of a crashing ocean wave, though it is unclear whether we are witnessing the end of the world’s final, apocalyptic storm.
TABLE V: Roberta Movie Generations based on 50% random context.
Table V shows that the generated texts have different levels of coherence and grammatical correctness. Overall, the summaries convey the main storylines smoothly and introduce the central characters at the right time; however, the narration is disrupted multiple times because of the hallucinations introduced by the model.
Many of the transitions in the summaries happened all of a sudden, which makes it harder to follow the plots. The second summary, for example, jumps from the childhood of the protagonist to his adult life. In addition, the generated text sometimes includes irrelevant details that draw attention away from the main storyline, such as the "Glock" and the slow movement of the desk in the fifth summary, which seem totally out of context. The model is sometimes limited in its ability to maintain long-term dependencies between characters in the plots; for example, in the fourth plot, the summary contradicts itself multiple times about the actions of Duke and Meta.
Nonetheless, the generations are mostly understandable at least. The main characters, settings, and conflicts are all present in the summaries. They are also easy to read, have clear language, and are similar to the language used in the original training set. The summaries are very different in style and tone, showing that the models can adapt to different genres.
CLM Generations These samples were generated by Bart large.
With montage scenes of daily Australia 1970s: desert main of a jam in the, being unusually adverse weather that appear only the aboriginal to recognize the significance of weather. The Australian legal aid, (Chamberlain) defence the circumstances by which, in that taxation the and personal lives unravel. He is plagued by bizarre dreams begins to connection to one of the accused (David Gulpilil) also feels connected increasingly strange phenomena city. His intensify along with the case, which he comes an aboriginal killing curse in which the victim about aboriginal and the as a parallel world of existence, Burton comes to weather bodes the end of the world. Overcoming the climaxes confrontation the lawyer and tribe’ subterranean sacred site. After the sewer, he on beach and stares. Screen filled by the shot of a ocean wave unclear if he is witnessing or in’ final, he stares at the ocean. The film ends with one of these "freak" Sydney an occurs among a of in the mysterious of the mysterious death of one of at coroner, death ruled homicide; men. Burton escapes to in the tunnel various relics. Overcoming, he escapes to the beach, where he stares to the ocean and stares at it.
TABLE VI: Bart Movie Generations based on 50% random context.
Table VI shows that overall, the quality of these writings is far below that of the MLM generations. The primary issue is the lack of coherence in specific passages as the text occasionally struggles to maintain a clear narrative flow, with abrupt transitions and unclear connections between each sequence. Most passages contain vague statements that leave room for interpretation; they seem more like a collection of random sentences lacking logical connections between them. Another critical problem is that the texts contain grammatical errors, such as incorrect verb tenses, subject-verb agreement issues, and awkward sentence structures. Additionally, the text sometimes lacks clarity and precision in its language, with the occasional appearance of random characters in the text.
IV-B3 Author Generations
Below are the generations of Roberta-large and Bart-large. It should be noted that some of the problems the models show with these generations were also present in the original dataset, so the training data has probably influenced the quality of the generated texts.
MLM Generations These samples were generated by Roberta large.
’hi, thank you so much for clarifying this, im so sorry for the inconvenience i am happy to take all the shifts for tomorrow, again thank you for your help Hi, good afternoon, i have been unable to upload my powerpoint presentation as an attachment as the file is not working i have attempted to upload it as an onedrive, however i am experiencing issues could i please upload the slides from my memory stick if any further attempts fail i will continue to continue to attempt to upload if you could let me know if this is working i would be very grateful kindest regards, Hi, i have estimated as an end date for tomorrow hope this is okay truly, Hi, i have attached pictures of my presentation, as have many others, good afternoon, i hope your well im sorry for the delayed reply, i was unable to hear back from you guys about my dbs, which has now been received i have scanned my dbs but they are much bigger than a4x5 printer normally scans so they will not show all the information on each side i was able to print out all the documents on hard drive i have attached the pdf containing the information required per the interview, there are also 2 pdfs containing my CV and email address thank you for your help, Hi, my lovelies, i hope youre all well i am so happy to inform you i have secured the role with i would not have got through the interview without all your support and help, i cant thank you enough for your help with my placement i hope you guys will be able to answer if i am missing anything in any of my questions, i can tell you that i will be completing my placement for 100 days, rather than a year anything which is very useful to carry around day i am currently searching for a diary what is training and is it something i should bringwear me as i havent been provided with a location yet, will routes also be required to attend training, or will it depend on the number of days i will be on a placement is there anything you can help me with before starting, any information i can get is much more beneficial to me i will be contacted via your website, once again thank you for your continuous,’
TABLE VII: Roberta Author Generations based on 50% random context.
Table VII shows that one significant issue throughout the texts is the lack of proper grammar, punctuation, and sentence structure. The model struggles to maintain coherence and clarity, often producing fragmented and poorly organized ideas. In addition the texts are full of repetitive phrases, lack of smooth transitions between ideas. However, the model does have the ability to capture different writing styles, such as casual conversations or more formal tones. Overall, the texts are understandable and do not show any significant errors.
Hi thank you so clarifying that im will bring the for once thank you, good afternoon, i am unable to my powerpoint presentation as attachment as the is too big to it as onedrive i am experiencing some issues with my memory, could i the presentation memory all other attempts fail i will however to you this is i be very grateful kindest regards, date placement i hope yours, hi have attached dbs thanks, Hi , I have attached my dbs Thank you for your support, Hi, I hope youre well, I have a queries the placement hope will be answer if thats ok to of my questions can tell you that i be for days over a will to carry day to day i am currently what training like should i been with location yet so routes would be helpful recommendations the prices on number days i will completing you guys before placement any you think relevant for to know i contacted this again thank you for my support, thank you enough Good afternoon, My name is and I am a student at , I am on placement for with I am so pleased to the role with i not gone through the application process as i have not received any support, Thank you enough, Hello, your have attached the dbs to my presentation Thanks, Good morning, my university about form has now i have as bigger than a4 only a4 show all each am to documents as copies 4 the required per there are also 2 attached containing address you thanks, hi, i hope Youre well im delayed i have been contacted by the university about my form, my university has now informed me that my form has been sent to for me to have a look at it thanks,
TABLE VIII: Bart Author Generations based on 50% random context.
CLM Generations These samples were generated by Bart large. In Table VIII, the texts cover various topics and styles, including personal communication, academic writing, and conversational transcripts. However, the overall quality of the generated content is poor and lacks coherence, proper grammar, and punctuation. As with the movie synopses, the model struggles to maintain clarity, generating fragmented sentences. The information presented is often confusing and contains random or irrelevant details. Overall, the models cannot maintain a clear focus and coherent flow throughout each text.
IV-C Downstream Evaluation Tasks
In this section, we evaluate the generated texts used in three downstream tasks: NER, Text Classification, and Authorship Verification. These results will suggest the relationship between the quality of the generations and their usefulness in a practical setting, assessing the potential for use as datasets for different NLP tasks. In Section III-D we explained how each downstream task was designed.
IV-C1 Downstream NER Task
Figure 11: Downstream NER Evaluation Results.
Figure 11 shows that while the models trained on the real data perform slightly better (orange line), the models trained on the generated data also perform well. The graph shows a similar behaviour as the quantitative metric graphs. It shows good scores for high ratios, and it descends, although less decisively, for lower contexts. In contrast with the quantitative metrics, however, synthetic data generated via stopwords and punctuation as context can train the models to a comparable performance to random masking. For MLM models, NER tokens provide enough context to train the model effectively.
What is remarkable about these results is that even when the ratio is at 100%, the f-1 sharply decreased. However, it’s still able to effectively train the model, achieving an f-1 of almost 75%. This contrasts with the low quantitative scores achieved by this ratio. This shows that to train NER models effectively, the quality of the generated samples has a limited impact on the performance.
IV-C2 Downstream Text Classification Task
Figure 12: Downstream Text Classification Evaluation Results.
The results show that the models trained on the generated data perform significantly differently based on the quality of the generated texts. The worst model had an f-1 score of 0.42, and the worst model was 0.36. Even more evidently than in the NER task, the quality of the text does not correlate with the performance of the model; as a matter of fact, the Roberta-large 100% masking ratio performs better than most other ratios and punctuation for t5-large performs even better than the original training data.
It should be noted that given the generation constraints, we could train the model only on 200 samples; therefore, the models don’t perform excellently in this task.
IV-C3 Downstream Author Verification Evaluation
Figure 13: Downstream Authorship Verification Evaluation Results.
The generations show performances comparable to the quantitative metrics for 3 out of the 4 models. This is a good sign which means that the models can leverage the context appropriately. For the MLM models, however, the performance decreases sharply at the lower end of the ratios (90 - 100%). T5-large, in contrast, shows no correlation between the number of matches and the amount of context given, meaning that the model is indifferent to the amount of context given.
What is peculiar is also the fact that punctuation alone obtains the highest performance of all for the CLM models (T5 & Bart); however, when looking at the actual generations, we realized that the models show clear hallucinations when given just punctuation, in fact all the generations are just ",,,,," leading the prediction model to predict correctly predict "Yes" (text 1 and the reconstructed text come from the same author) but with low confidence.
With regard to stopwords and stopwords&punctuation, they show fairly good performance, placing in the middle of the graph. Punctation does not give enough context to the MLM models and shows very low performance, which is shown as the lowest in the Figure.
IV-D Discussion
The quantitative and qualitative evaluations of this study, as well as the performance shown in the downstream tasks, provide valuable insights into the performance of the different language models and which are the best masking approaches for text generation.
Our main finding is that 1) MLM consistently outperformed CLM in text generation across all three datasets. The quantitative metrics indeed showed that the texts generated by MLM models were more aligned with the original reference texts. The qualitative human evaluation showed that MLM generations had better coherence and were more grammatically correct.
2) Incorporating domain-specific knowledge in the generation models did not improve performance. This was shown by the scores obtained with BiomedNLP-PubMedBERT and SciFive-Large-Pubmed_PMC, two models focused on the biomedical domain, did not achieved better scores than the general models in any of the quantitative metrics.
3) We found no strong correlation between the quality of the generated texts and the performance in the downstream tasks. While most models trained on the generations of higher quality did perform well in the downstream tasks, some of the models trained on lower-quality data (e.g., 100% masking ratio) still achieved good results. We believe that this is a very interesting discovery, and it suggests that downstream tasks are more resilient to the noise that the generative models add with their generations. It is also possible that the synthetic generations make the models more robust since they introduce significant variability in the training sets.
4) As expected, random masking with lower masking ratios (preserving more context) led to higher-quality generations. However, even with limited context, such as NER or stopwords_punctuation masking, MLM models could generate texts of good overall quality.
V Conclusions and Future Work
In this paper, we compared CLM and MLM text generation across medical, movie synopsis, and authorship verification datasets. We explored different masking approaches to determine the most effective ones. We then assessed the quality of the generations first with quantitative and qualitative evaluations, and then we assessed their usability by running three downstream tasks. Our results established that MLM can generate better text than CLM.
V-A Limitations
Our findings provide valuable insights into the performance of MLM and CLM approaches for text generation. However, in this section, we want to acknowledge some of the limitations that may impact the generalisability of our results.
Computational Efficiency: The iterative process of generating text using MLM models, particularly with high masking ratios, can be computationally expensive compared to autoregressive CLM models. Each iteration requires feeding the entire masked sequence through the model, which can be time-consuming for longer texts. This is because The current libraries are not optimized for generating texts using MLM models.
Quantitative Evaluation Metrics: The quantitative evaluation relied upon n-gram-based metrics because they provide a valuable indication of text quality; however, it should be noted that they were initially designed for tasks such as machine translation and text summarisation. As such, these measures may not fully align with the kind of text that we are looking for. In addition, it should be the MLM models had an advantage in this evaluation because part of the real tokens were already in the right place. Therefore, it is possible that a higher score does not correspond to better text quality in this case.
Length of the Texts: The BERT family of models, including those we employed, has a maximum input sequence length of 512. This may limit the application of MLM-based text generation in domains that usually involve longer texts.
Size of the Models: Due to computational resources and the fact that encoder-only models are generally small, we had to restrict our study to models that have hundreds of millions of parameters. Therefore, their performance may be limited compared to larger state-of-the-art models with billions of parameters, and further research should focus on that instead.
V-B Future Work
Our comparison of MLM and CLM for text generation opened up multiple directions for future research. In this section, we outline four potential developments that could expand on our paper.
Creative Writing: Future work could involve our method to generate more unstructured text generated in tasks such as creative writing. This will help us understand whether our approach can not only enhance the quality of the generations but also their creativity.
Iterative Refinement: Iterative refinement [32] involves gradually improving the initial generations over multiple iterations until good text quality is achieved. When applied to MLM, this process could entail identifying and masking problematic tokens or text sequences and then regenerating those parts to improve the quality of the final output.
More Downstream Tasks: This paper showed that texts of higher quality do not necessarily lead to higher performance in the downstream tasks. In this work, we focused on NER, text classification and authorship verification; however, future work should include a broader range of NLP tasks, such as sentiment analysis, question answering and POS tagging. By evaluating the performance of models trained on MLM-generated data across a diverse set of tasks, we will be able to show more definitively the versatility and robustness of this method.
On Medical Data: MLM models showed the best performance compared to CLM in the medical dataset so future research could investigate the application of MLM for data augmentation and synthetic data generation in the biomedical domain further. This is, in fact, the domain that needs this kind of data the most because it is generally hard to access due to privacy concerns [7]. Future research could involve more masking strategies and text types, as well as evaluating the generated data on a range of medical downstream tasks. Another critical aspect to add to this research would be to collaborate with actual clinicians to assess the plausibility of the generated text along with its quality; this could provide us valuable insights into the possible use of this approach in clinical settings.